OpenStudy (anonymous):

Find the equation of a polynomial that has the roots 3,2, and -1 and goes through the point 1,1

OpenStudy (anonymous):

reverse engineer the quadratic equation to fit specs

OpenStudy (dumbcow):

y = a(x-3)(x-2)(x+1)

OpenStudy (anonymous):

@ dumbcow What do I do with "a"?

OpenStudy (dumbcow):

its a constant we solve for it by putting in the x and y values from the point (1,1) 1=a(1-3)(1-2)(1+1)

OpenStudy (anonymous):

Oh I see. I thought I could just add one on to the equation y=(X-3).... This makes sense. Thank you.

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

@ dumbcow. Just finished the problem... It works!

OpenStudy (anonymous):

u could also use -b/a = x1+x2 and c/a = x1x2 and solve for a and b or c then use the fact that f[1]=1 to find c or b